OpenStudy (anonymous):

1. Find X 2. Find Y A) 8 B) 12 C) 15 D) 7(squareroot)3 E) 4 AB) 2 AC) 15(squareroot)3 / 2 AD) 7(squareroot)3 / 3 AE) 4(squareroot)3 BC) 16 BD) 8(squareroot) 3 BE) 30(squareroot) 3 CD) 15/2 CE) 16(squareroot) 3 DE) 14 ABC) 14(squareroot) 3 / 3

OpenStudy (anonymous):



myininaya (myininaya):

y=2x x^2+3^2=y^2 but y=2x x^2+9=(2x)^2 x^2+9=4x^2 subtract x^2 on both sides 9=3x^2 so x^2=3 now this you help you finish